Solve by substitution or elimination:
[x+Y=13
[2x-3y=1
1
Expert's answer
2012-10-15T11:20:18-0400
x+y=13 => y=13-x&
2x-3y=1

2x-3(13-x)=1
2x-39+3x=1
5x=40
x=8& => y=13-x=13-8=5